OpenStudy (anonymous):

Leo bought a bulldozer for $63,103. The value of the bulldozer depreciated at a constant rate per year. The table below shows the value of the bulldozer after the first and second years: Year 1 2 Value(in dollars) 58054.76 53410.38 Which function best represents the value of the bulldozer after t years? a) f(t) = 58,054.76(0.92)t b) f(t) = 63,103(0.08)t c) f(t) = 63,103(0.92)t d) f(t) = 58,054.76(0.08)t

OpenStudy (danjs):

\[f(t) = A*b^t\] They give you two points, you use them to find the constants A and b

OpenStudy (danjs):

with the initial value t=0 being 63103

OpenStudy (danjs):

at time t=0 f(t) = A*1 \[b^0 = 1\]

OpenStudy (anonymous):

so what answer would it be?

OpenStudy (danjs):

Using t=0 f(0)=A (t,f(t)) = (0 , 63103) f(0) = A = 63103 Soyou have the value for A

OpenStudy (danjs):

\[f(t) = 63103*b^t\] use one of the points (t , f(t)) in the equation, and solve for b

OpenStudy (danjs):

t= 1 , f(t) = 58054.76 58054.76 = 63103*b^1 solve for b

OpenStudy (danjs):

Use that b, in \[f(t) = 63103*b^t\] that is like 99% of the work, i cant do more

OpenStudy (anonymous):

b=0.92?

OpenStudy (danjs):

right

OpenStudy (anonymous):

yaay

OpenStudy (danjs):

i hope you see the process to get A and b though, in f(t) = A*b^t

OpenStudy (danjs):

A = initial value when t=0

OpenStudy (danjs):

then use any of the points they gave (t,f(t)) and calculate b

OpenStudy (anonymous):

would the answer be b?

OpenStudy (danjs):

you told me what the value of b is in b=0.92 A=initial value =63103 f(t) =A*b^t

OpenStudy (anonymous):

i solve for t?

OpenStudy (danjs):

no, where you see A, you put 63103 where you see b, you put 0.92 done

OpenStudy (anonymous):

oh i got 58054.76

OpenStudy (danjs):

\[f(t) = 63103*(0.92)^t\] what are you solving, how did you get 58054.76

OpenStudy (anonymous):

oh i multiplied 36103 * 0.92

OpenStudy (danjs):

you cant do that, 0.92 is raised to a power

OpenStudy (danjs):

remember, powers are calculated before multiplication

OpenStudy (anonymous):

well whats the power

OpenStudy (danjs):

t

OpenStudy (danjs):

for any time t you choose, the function f(t) will give you a value of the bulldozer at that time

OpenStudy (anonymous):

oh ok so its c

OpenStudy (danjs):

The question is looking for the general function f(t) for the depretiation of the bulldozer f(t) = 63103 * 0.92^t

OpenStudy (danjs):

The general exponential function looks like the form f(t) = A*b^t did you understand how the data from the problem was used to figure out A = 63103 and b=0.92?